OpenStudy (anonymous):

What is the vertex for the function? (You will want to find the axis of symmetry first.) Write as an ordered pair, leaving no spaces, such as (1,-4). y = -3x2 + 6x

OpenStudy (jdoe0001):

hmm for a quadratic equation like that \(\bf \Large ax^2+bx+c\) , which is just a parabola, you can find the vertex at the coordinates at \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (jdoe0001):

notiice that \(\bf y=-3x^2+6x \implies y=-3x^2+6x+0\)

OpenStudy (anonymous):

Okay, I see that.

OpenStudy (jdoe0001):

\(\bf y = -3x^2 + 6x+0\\ \quad \\ \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\implies \left(-\cfrac{(6)}{2(-3)}\quad ,\quad (0)-\cfrac{(6)^2}{4(-3)}\right)\)

OpenStudy (anonymous):

Okay, I'm sorry I'm not too good in math. Cam you explain this a bit more for me?

OpenStudy (jdoe0001):

well.... is simple.... just evaluate the fractions

OpenStudy (anonymous):

okay, so (6).2(-3), and 6^2/4(-3)?

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} y = -3x^2 + 6x+0\\ \quad \\ \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\implies &\left(-\cfrac{(6)}{2(-3)}\quad ,\quad (0)-\cfrac{(6)^2}{4(-3)}\right)\\ &\qquad x\qquad \qquad \qquad \qquad y \end{array}\)

OpenStudy (anonymous):

Okay, awesome, thanks!

OpenStudy (jdoe0001):

yw